OpenStudy (anonymous):

How do I solve this? Please show step by step 1/2(5/6+2/7x)=19/42

OpenStudy (zzr0ck3r):

\[\frac{1}{2}(\frac{5}{6}+\frac{2}{7}x)=\frac{19}{42}\]

OpenStudy (zzr0ck3r):

distribute the 1/2 threw

OpenStudy (zzr0ck3r):

or multiply both sides by 2...either way

OpenStudy (anonymous):

Ok